# The Landscape of GPU-Centric Communication

### v2

@@ 4.3.1. NVSHMEM
NVSHMEM is NVIDIA’s implementation of the OpenSHMEM specification for CUDA devices. NVSHMEM is a Partitioned Global Adress Space (PGAS) library that provides efficient one-sided put / get APIs for processes to access remote data objects. NVSHMEM supports point-to-point and collective communication between GPUs both within and across nodes (NVIDIA, 2023i).
@@
A notable attribute of NVSHMEM is that it offers both host-side and device-side APIs. The host-side APIs expose an optional stream argument that can be used to implement communication-computation overlap. For certain calls, the GPU-side variants provide the calls in three granularities: thread, thread block, and warp. The thread variant means that the call should be performed by a single device thread and will be executed by that thread. The thread block and warp variants use multiple threads to execute the communication call cooperatively. These variants should be called by all threads in the corresponding thread block or warp. A previous performance comparison between the host-side and device-side APIs found negligible differences in performance between the two, with host-side APIs slightly outperforming the device-side variants (Groves et al., 2020). This study was conducted using an early version of NVSHMEM (0.3.0), which has since seen improvements in GPU-side API performance.
@@
As of version 2.7.0, NVSHMEM introduced the Infiniband GPUDirect Async (IBGDA) transport built on top of GPUDirect Async (NVIDIA, 2023j). The IBGDA transport allows GPUs to issue inter-node communication directly to the NIC, bypassing the CPU entirely. Without IBGDA, device-side inter-node communication calls are performed through a proxy thread on the CPU that triggers the corresponding NIC operations. This proxy thread consumes CPU resources and creates a bottleneck in achieving peak NIC throughput for fine-grained transfers (Pak Markthub and Howell, 2022). NVSHMEM with IBGDA support, combined with persistent kernels, enables the complete transfer of both data and control paths to the GPU and marks a significant shift towards fully autonomous multi-GPU execution. However, as discussed in Section 3.2.3, GPUDirect RDMA only enforces GPU-NIC memory consistency across kernel boundaries. This inherent reliance on the CPU for memory consistency is a potential obstacle toward truly autonomous multi-GPU execution. One workaround is using a callback mechanism whereby the persistent kernel signals the CPU to perform a consistency-enforcing API call (i.e., cudaDeviceFlushGPUDirectRDMAWrites()). The efficacy of this solution is unclear and warrants further investigation. Enforcing GPU-NIC memory from inside the kernel is supported by ROC_SHMEM, which we discuss next.
@@ 4.3.2. ROC_SHMEM
ROC_SHMEM is AMD’s implementation of the OpenSHMEM specification for AMD GPUs. ROC_SHMEM offers two communication backends. The first, known as GPU-IB, implements Infiniband on the GPU, similar to NVSHMEM’s IBGDA transport. The second, called Reverse Offload (RO), uses host-side proxy threads and offloads communication to the CPU. GPU-IB is the default backend and offers the best performance (AMD, 2023e). ROC_SHMEM works almost identically to NVSHMEM and offers analogous APIs. However, there are several significant differences. First, as mentioned in the previous section, NVSHMEM runs into GPU-NIC memory consistency problems when intra-node communication is issued from persistent kernels. ROC_SHMEM, on the other hand, explicitly addresses this issue and guarantees correctness when persistent kernels are being used. Hamidouche et al. analyze the GPU-NIC memory mismatch stemming from the GPU’s relaxed memory model and propose changes integrated into ROC_SHMEM (Hamidouche and LeBeane, 2020). This means that ROC_SHMEM provides completely CPU-free communication mechanism that can move the entire flow of multi-GPU execution to the device.
@@
Second, ROC_SHMEM uses GPU shared memory (local data store (LDS) in AMD parlance) to store network state for faster access. To the best of our knowledge, this optimization is not implemented in NVSHMEM. While this is most likely beneficial for execution time, the increased shared memory message could limit occupancy and negatively impact performance (Punniyamurthy et al., 2023a; AMD, [n. d.]).
@@
Third, prior versions of ROC_SHMEM required allocating symmetric buffers as uncacheable in order to prevent stale data from being communicated. However, as AMD recently introduced intra-kernel cache flush instructions, the data can be flushed before initiating the network transaction, allowing the data to be cached. No such instructions are provided by NVIDIA, meaning that NVSHMEM buffers are likely allocated as uncacheable (Punniyamurthy et al., 2023a).
4.3.3. Applications
In recent years, NVSHMEM has been integrated as a communication backend into multiple runtimes. PETSc implemented PetscSF, a scalable communication layer based on NVSHMEM, to complement their MPI-based approach, which did not work well with CUDA stream semantics and prevented kernel launch pipelining (Zhang et al., 2021). Kokkos Remote Spaces, adding distributed memory support to the Kokkos programming model, uses NVSHMEM as a communication backend (Ciesko, 2023; Trott, 2018). An NVSHMEM implementation of the Kokkos Conjugate Gradient Solver outperforms the CUDA-aware MPI implementation while significantly reducing the code base size (Maruyama et al., 2020). Choi et al. use persistent kernels and NVSHMEM to implement CharminG, a fully GPU-resident runtime system inspired by Charm++ (Choi et al., 2021b). The Livermore Big Artificial Neural Network (LBANN) implements a spatial-parallel convolution using NVSHMEM that outperforms MPI and Aluminium implementations (Maruyama et al., 2020). QUDA, a library for lattice QCD computations, has used NVSHMEM and persistent kernels for improved strong scaling of Dirac operators (lattice, 2023; Wagner, 2020).
NVSHMEM has also been used outside of runtime-based approaches to achieve performance improvements. Chu et al. combine NVSHMEM with persistent kernels to implement a state-of-the-art GPU-based key-value store (Chu et al., 2019). Xie et al. use NVSHMEM to implement a single-node multi-GPU sparse triangular solver (SpTRSV) that achieves good performance scalability compared to a UVM-based design (XIE et al., 2021). Ding et al. combine persistent kernels with NVSHMEM for impressive performance in a sparse triangular solver (SpTRSV) on single- and multi-node setups. Atos implements both persistent and discrete kernels with NVSHMEM-based communication to achieve state-of-the-art performance on multi-GPU BFS within and across nodes (Chen et al., 2022a). Wang et al. propose MGG, a system design that accelerates Graph Neural Networks (GNNs) on multi-GPU systems using a GPU-centric software communication-computation pipeline with NVSHMEM for fine-grained communication (Wang et al., 2023). Ismayilov et al. use persistent kernels and device-side NVSHMEM to implement fully GPU-side Jacobi 2D/3D and CG solvers that outperform CPU-controlled baselines. They reserve some thread blocks for communication while others handle computation, a technique they call thread block specialization, to achieve explicit device-side communication-computation overlap (Ismayilov et al., 2023). Punniyamurthy et al. use ROC_SHMEM and persistent kernels to overlap embedding operations with collective communication in deep learning recommendation models (Punniyamurthy et al., 2023a).

### v3

@@ 4.3.1. NVSHMEM
NVSHMEM is NVIDIA’s implementation of the OpenSHMEM specification for CUDA devices. NVSHMEM is a Partitioned Global Adress Space (PGAS) library that provides efficient one-sided put / get APIs for processes to access remote data objects. NVSHMEM supports point-to-point and collective communication between GPUs both within and across nodes (NVSHMEM).
@@
A notable attribute of NVSHMEM is that it offers both host-side and device-side APIs. The host-side APIs expose an optional stream argument that can be used to implement communication-computation overlap. For certain calls, the GPU-side variants provide the calls in three granularities: thread, thread block, and warp. The thread variant means that the call should be performed by a single device thread and will be executed by that thread. The thread block and warp variants use multiple threads to execute the communication call cooperatively. These variants should be called by all threads in the corresponding thread block or warp. A previous performance comparison between the host-side and device-side APIs found negligible differences in performance between the two, with host-side APIs slightly outperforming the device-side variants (Performance-Trade-offs-in-GPU-Communication-A-Study-of-Host-and-Device-Initiated-Approaches). This study was conducted using an early version of NVSHMEM (0.3.0), which has since seen improvements in GPU-side API performance.
@@
As of version 2.7.0, NVSHMEM introduced the Infiniband GPUDirect Async (IBGDA) transport built on top of GPUDirect Async (NVSHMEM-2.7.0-Release-Notes). The IBGDA transport allows GPUs to issue inter-node communication directly to the NIC, bypassing the CPU entirely. Without IBGDA, device-side inter-node communication calls are performed through a proxy thread on the CPU that triggers the corresponding NIC operations. This proxy thread consumes CPU resources and creates a bottleneck in achieving peak NIC throughput for fine-grained transfers (Blog-Improving-Network-Performance-of-HPC-Systems-Using-NVIDIA-Magnum-IO-NVSHMEM-and-GPUDirect-Async). NVSHMEM with IBGDA support, combined with persistent kernels, enables the complete transfer of both data and control paths to the GPU and marks a significant shift towards fully autonomous multi-GPU execution. However, as discussed in Section 3.2.3, GPUDirect RDMA only enforces GPU-NIC memory consistency across kernel boundaries. This inherent reliance on the CPU for memory consistency is a potential obstacle toward truly autonomous multi-GPU execution. One workaround is using a callback mechanism whereby the persistent kernel signals the CPU to perform a consistency-enforcing API call (i.e., cudaDeviceFlushGPUDirectRDMAWrites()). The efficacy of this solution is unclear and warrants further investigation. Enforcing GPU-NIC memory from inside the kernel is supported by ROC_SHMEM, which we discuss in the next section.
In recent years, NVSHMEM has been integrated as a communication backend into multiple runtimes. PETSc implemented PetscSF, a scalable communication layer based on NVSHMEM, to complement their MPI-based approach, which did not work well with CUDA stream semantics and prevented kernel launch pipelining (PetscSF-Scalable-Communication-Layer). Kokkos Remote Spaces, adding distributed memory support to the Kokkos programming model, uses NVSHMEM as a communication backend (Kokkos-Remote-Spaces-Repository; Talk-Early-Experience-with-NVSHMEM-Extending-the-Kokkos-Programming-Model-with-PGAS-Semantics). An NVSHMEM implementation of the Kokkos Conjugate Gradient Solver outperforms the CUDA-aware MPI implementation while significantly reducing the code base size (CUDA-Blog-Scaling-Scientific-Computing-with-NVSHMEM). Choi et al. use persistent kernels and NVSHMEM to implement CharminG, a fully GPU-resident runtime system inspired by Charm++ (CharminG-A-Scalable-GPU-Resident-Runtime-System). The Livermore Big Artificial Neural Network (LBANN) implements a spatial-parallel convolution using NVSHMEM that outperforms MPI and Aluminium implementations (CUDA-Blog-Scaling-Scientific-Computing-with-NVSHMEM). QUDA, a library for lattice QCD computations, has used NVSHMEM and persistent kernels for improved strong scaling of Dirac operators (QUDA-Repository; Talk-NVSHMEM-Overcoming-Latency-Barriers).
NVSHMEM has also been used outside of runtime-based approaches to achieve performance improvements. Chu et al. combine NVSHMEM with persistent kernels to implement a state-of-the-art GPU-based key-value store (Designing-High-Performance-In-Memory-Key-Value-Operations-with-Persistent-GPU-Kernels-and-OpenSHMEM). Xie et al. use NVSHMEM to implement a single-node multi-GPU sparse triangular solver (SpTRSV) that achieves good performance scalability compared to a UVM-based design (Fast-and-Scalable-Sparse-Triangular-Solver-for-Multi-GPU-Based-HPC-Architectures). Ding et al. combine persistent kernels with NVSHMEM for impressive performance in a sparse triangular solver (SpTRSV) on single- and multi-node setups. Atos implements both persistent and discrete kernels with NVSHMEM-based communication to achieve state-of-the-art performance on multi-GPU BFS within and across nodes (Atos-Vyse-Scalable-Irregular-Parallelism-with-GPUs-Getting-CPUs-Out-of-the-Way). Wang et al. propose MGG, a system design that accelerates Graph Neural Networks (GNNs) on multi-GPU systems using a GPU-centric software communication-computation pipeline with NVSHMEM for fine-grained communication (MGG-Accelerating-Graph-Neural-Network-with-Fine-grained-intra-kernel-Communication-Computation-Pipelining-on-Multi-GPU-Platforms). Ismayilov et al. use persistent kernels and device-side NVSHMEM to implement fully GPU-side Jacobi 2D/3D and CG solvers that outperform CPU-controlled baselines. They reserve some thread blocks for communication while others handle computation, a technique they call thread block specialization, to achieve explicit device-side communication-computation overlap (Multi-GPU-Communication-Schemes-for-Iterative-Solvers-When-CPUs-Are-Not-in-Charge). Punniyamurthy et al. use ROC_SHMEM and persistent kernels to overlap embedding operations with collective communication in deep learning recommendation models (GPU-initiated-Fine-grained-Overlap-of-Collective-Communication-with-Computation).
@@ 4.3.2. ROC_SHMEM
ROC_SHMEM is AMD’s implementation of the OpenSHMEM specification for AMD GPUs. ROC_SHMEM offers two communication backends. The first, known as GPU-IB, implements Infiniband on the GPU, similar to NVSHMEM’s IBGDA transport. The second, called Reverse Offload (RO), uses host-side proxy threads and offloads communication to the CPU. GPU-IB is the default backend and offers the best performance (ROCSHMEM). ROC_SHMEM works almost identically to NVSHMEM and offers analogous APIs. However, there are several significant differences. First, as mentioned in the previous section, NVSHMEM runs into GPU-NIC memory consistency problems when intra-node communication is issued from persistent kernels. ROC_SHMEM, on the other hand, explicitly addresses this issue and guarantees correctness when persistent kernels are being used. Hamidouche et al. analyze the GPU-NIC memory mismatch stemming from the GPU’s relaxed memory model and propose changes integrated into ROC_SHMEM (GIO-GPU-Initiated-OpenSHMEM-Correct-and-Efficient-Intra-Kernel-Networking-for-dGPUs). This means that ROC_SHMEM provides completely CPU-free communication mechanism that can move the entire flow of multi-GPU execution to the device.
@@
Second, ROC_SHMEM uses GPU shared memory (local data store (LDS) in AMD parlance) to store network state for faster access. To the best of our knowledge, this optimization is not implemented in NVSHMEM. While this is most likely beneficial for execution time, the increased shared memory message could limit occupancy and negatively impact performance (GPU-initiated-Fine-grained-Overlap-of-Collective-Communication-with-Computation; AMD-Instinct-MI200-Instruction-Set-Architecture).
@@
Third, prior versions of ROC_SHMEM required allocating symmetric buffers as uncacheable in order to prevent stale data from being communicated. However, as AMD recently introduced intra-kernel cache flush instructions, the data can be flushed before initiating the network transaction, allowing the data to be cached. No such instructions are provided by NVIDIA, meaning that NVSHMEM buffers are likely allocated as uncacheable (GPU-initiated-Fine-grained-Overlap-of-Collective-Communication-with-Computation).
4.3.3. Intel SHMEM
Intel recently introduced Intel SHMEM (IntelShmem), the first GPU-aware implementation of OpenSHMEM for Intel GPUs. It allows SHMEM routines to operate directly on GPU memory and supports GPU-initiated communication by embedding SHMEM calls inside SYCL kernels. Intel SHMEM integrates with the SYCL programming model, offering a portable C++ interface for heterogeneous systems. In contrast, existing solutions such as NVSHMEM and ROC_SHMEM provide similar capabilities but are tied to vendor-specific ecosystems.
To maintain compatibility with the OpenSHMEM 1.5 specification, Intel SHMEM supports both device- and host-side APIs for point-to-point operations, collective operations via the teams API, and SYCL-specific extensions for work-group and sub-group communication. Inter-node communication is handled through a host proxy thread, which forwards GPU-initiated operations to a standard OpenSHMEM backend. In other words, GPU issues SHMEM operations inside a SYCL kernel. The GPU writes work queue entries into a memory region accessible by both GPU and host. The host thread running on CPU detects the GPU-initiated RMA operation. The host executes actual RDMA using a standard OpenSHMEM library. The current implementation relies on Sandia OpenSHMEM (SOS), leveraging its robust support for OFI/libfabric transports and its capability to maintain a symmetric heap in GPU memory. This layered design enables Intel SHMEM to provide functional equivalence with NVSHMEM and ROC_SHMEM while integrating cleanly with Intel’s broader oneAPI and SYCL ecosystem.
For intra-node communication, Intel SHMEM can perform GPU-to-GPU data movement directly without involving the host CPU. This is achieved by allowing the GPU to issue SHMEM operations from within SYCL kernels, using Intel’s Level Zero–based backend to translate these operations into: (1) direct GPU load/store transfers, when GPUs share a unified memory fabric, or (2) GPU copy-engine transfers, which bypass the host CPU and use dedicated DMA engines on the GPU.
4.4. Comparison and Discussion of User-Level Libraries
While GPU-aware MPI, GPUCCL, and GPUSHMEM all offer mechanisms for programming multi-GPU systems, significant differences exist in their semantics and performance characteristics. Key distinctions include stream support, API location, programming and performance.
4.4.1. Streaming Support
GPUs operate on the concept of streams which are command queues that guarantee ordering among GPU operations. The GPU scheduler ensures that kernels and other operations launched on a stream execute in the order they were enqueued and do so with correct data dependencies. Since kernel launches are asynchronous and do not block the host, GPU runtimes can pipeline kernel launches and overlap the launch latencies behind kernel execution. The semantic mismatch between the MPI and GPU models is that MPI has no awareness of GPU streams. As a result, it is not possible to enqueue an MPI call on a given GPU stream or for a GPU stream to wait on the completion of a pending MPI routine. By implication, interlacing MPI calls with GPU kernels will require host-blocking synchronizations in order to maintain data correctness. For example, before initiating an MPI send, the programmer has to block the host to synchronize all streams which operate on the send buffer. Similarly, waiting on completion of pending MPI communication will also require host-blocking synchronization. In either case, these forced synchronizations impair kernel launch pipelining, prevent opportunities for overlap and force the programmer into alternating bulk phases of communication and computation (Aluminum-An-Asynchronous-GPU-Aware-Communication-Library-Optimized-for-Large-Scale-Training-of-Deep-Neural-Networks-on-HPC-Systems; PetscSF-Scalable-Communication-Layer; CPU-and-GPU-initiated-Communication-Strategies-for-CG; uniconn).
We see two possible non-mutually exclusive paths for resolving the semantic mismatch. The first is making MPI runtimes stream-aware by adding an explicit stream parameter to MPI routines. This would solve the issue of impaired kernel launch pipelining and allow MPI calls to seamlessly integrate into GPU runtimes. The second is providing the option of device-initiated MPI calls. This would reduce the programmer burden of juggling two distinct programming models and, additionally, provide implicit communication-computation overlap. Both directions have been explored in the literature on a limited scale. The FLAT compiler automatically converts device-side MPI calls to their host-side equivalents (Flat-A-GPU-Programming-Framework-to-Provide-Embedded-MPI). dCUDA implements device-side operations with MPI semantics but uses CPU helper threads for the actual communication. They rely on the GPU’s inherent memory latency hiding capabilities to implicitly overlap communication with computation, ultimately outperforming a GPU-aware MPI baseline (dCUDA-Hardware-Supported-Overlap-of-Computation-and-Communication). Namashivayam et al. explore new communication schemes to introduce GPU stream-awareness in MPI. They use the triggered operations feature on HPE Slingshot 11 interconnect, allowing the CPU to enqueue communication and synchronization operations to the NIC, which the GPU can then trigger. This reduces CPU involvement in the critical path and eliminates expensive synchronizations. While inter-node experiments show some performance benefits, the proposed scheme falters in intra-node setups as progress threads need to be used to emulate deferred execution semantics (Exploring-GPU-Stream-Aware-Message-Passing-using-Triggered-Operations). A follow-up work eliminates progress threads for intra-node communication, opting to use P2P Direct Load Store-based GPU kernels and GPU IPC-based mechanisms instead. The evaluation shows performance improvements over stream-oblivious MPI baselines (Exploring-Fully-Offloaded-GPU-Stream-Aware-Message-Passing).
More recently, MPIX streams (10.1145/3555819.3555820) allow an application to explicitly map its GPU execution stream context and pass it to the MPI library. This enables the MPI implementation to operate directly on GPU streams, thereby eliminating unnecessary synchronization overhead and improving the performance of GPU-to-GPU communication. However, none of these solutions have yet been incorporated into the official MPI standard.
4.4.2. Host vs Device-side API
Both GPU-aware MPI and GPUCCL utilize host-side APIs, mandating that communication routines and their parameters (like size and data type) are expressed on the host CPU using host arguments. GPUSHMEM offers both host and device-side APIs, enabling communication to be invoked directly from a GPU device kernel with input arguments defined as device variables. This device-centric approach allows GPUSHMEM to better handle low-latency and dynamic communication patterns. Data transfers are initiated within the computation kernel and sent straight to the network, minimizing latency and making the approach ideal for fine-grained communication.
However, this device-side invocation introduces programming and performance measurement challenges. Mixing computation and communication within a single kernel complicates the programming model, and since current profiling tools only operate at the kernel granularity, performance profiling is non-trivial. Moreover, especially when combined with persistent kernels, the technique causes resource contention as streaming processors and threads must be shared between the two. When employing techniques like dividing thread blocks (TBs) for overlap, synchronizing the dedicated communication TBs versus the computation TBs becomes difficult. While Thread Block Clusters (TBCs) can potentially mitigate this synchronization issue, their programming further complicates the overall programming complexity.
Mimicking NVSHMEM, NVIDIA NCCL 2.28 (NCCL-2.28-release-notes) introduced a new device-side API to enable fusing communication and compute. This was a significant departure from earlier versions where all NCCL operations were host-initiated. The new API allows GPU kernels to directly initiate data movement, and its use requires setting up data buffers with symmetric memory windows to facilitate direct GPU-to-GPU communication.
1//Host-side API
2if (rank == 0) { // process 0
3 for (int j = 0; j < window_size; ++j) {
4 MPI_Isend(send_buf, message_size, MPI_FLOAT, 1, 100, MPI_COMM_WORLD, send_request + j);
5 }
6 MPI_Waitall(window_size, send_request, reqstat);
7 MPI_Recv(recv_buf, 1, MPI_FLOAT, 1, 101, MPI_COMM_WORLD, reqstat);
8} else { // process 1
9 for (int j = 0; j < window_size; ++j) {
10 MPI_Irecv(recv_buf, message_size, MPI_FLOAT, 0, 100, MPI_COMM_WORLD, recv_request + j);
11 }
12 MPI_Waitall(window_size, recv_request, reqstat);
13 MPI_Send(send_buf, 1, MPI_FLOAT, 0, 101, MPI_COMM_WORLD);
14}
Listing 1: Simple one-way bandwidth benchmark with GPU-aware MPI
4.4.3. Programming Examples
Listings 1, 2, 3 present simplified one-way bandwidth benchmarks using GPU-aware MPI, NCCL, and device-side NVSHMEM. These examples highlight the semantic differences among the three communication libraries. While all three take a buffer pointer and size, MPI and NCCL rely on a two-sided communication model with synchronous send/receive semantics. In contrast, NVSHMEM employs a one-sided model: put/get operations are asynchronous with respect to the remote GPU. This distinction is visible in Listings 1 and 2 versus Listing 3, where NVSHMEM requires the sender to specify the receiver’s buffer address directly.
1//Host-side API
2if (rank == 0) { // process 0
3 ncclGroupStart();
4 for (int j = 0; j < window_size; ++j) {
5 ncclSend(send_buf, message_size, ncclFloat, 1, comm, stream);
6 }
7 ncclGroupEnd();
8 ncclRecv(recv_buf, 1, ncclFloat, 1, comm, stream);
9} else { // process 1
10 ncclGroupStart();
11 for (int j = 0; j < window_size; ++j) {
12 ncclRecv(recv_buf, message_size, ncclFloat, 0, comm, stream);
13 }
14 ncclGroupEnd();
15 ncclSend(send_buf, 1, ncclFloat, 0, comm, stream);
16}
Listing 2: Simple one-way bandwidth benchmark with NCCL
1//Device-side API
2__global__ void comm_kernel_send(...) {
3 nvshmemx_putmem_signal_nbi_block(recv_buf, send_buf, nx, signal_buf + blockIdx.x, 1, NVSHMEM_SIGNAL_ADD, 1);
4 grid.sync();
5 if (blockIdx.x == 0 && threadIdx.x == 0)
6 nvshmem_signal_wait_until(signal_buf, NVSHMEM_CMP_GE, i + 1);
7}
8
9__global__ void comm_kernel_recv(...) { // process 1
10 if (threadIdx.x == 0)
11 nvshmem_signal_wait_until(signal_buf + blockIdx.x, NVSHMEM_CMP_EQ, i + 1);
12 grid.sync();
13 if (blockIdx.x == 0)
14 nvshmemx_putmem_signal_nbi_block(recv_buf, send_buf, 1 * sizeof(real), signal_buf, 1, NVSHMEM_SIGNAL_ADD, 0);
15}
16//Host-side API
17if (rank == 0) {// process 0
18 nvshmemx_collective_launch(comm_kernel_send, dim3(window_size), dim3(1024), kernelArgs, 0, stream);
19} else { // process 1
20 nvshmemx_collective_launch(comm_kernel_recv, dim3(window_size), dim3(1), kernelArgs, 0, stream);
21}
Listing 3: Simple one-way bandwidth benchmark with Device-Side GPUSHMEM
MPI does not expose a stream parameter as discussed in Section 4.4.1, whereas NCCL and NVSHMEM operations are explicitly tied to a CUDA stream. NCCL also supports grouping multiple operations between groupStart/groupEnd to amortize launch overhead. Because MPI lacks stream semantics, MPI programs must explicitly synchronize CPU and GPU progress e.g., using Waitall to ensure completion. NVSHMEM kernels must be launched using a specialized collective-launch routine (e.g., nvshmemx_collective_launch), which also constrains the number of thread blocks due to the lack of GPU preemption.
Supporting multiple communication libraries often forces developers to reimplement communication backends, reducing productivity. To address this, Sagbili et al. proposed a unified communication interface that consolidates these models under a single API and demonstrated performance comparable to naive implementations (uniconn).
4.4.4. Performance
Several works compare the performance of different user-level libraries, either on real applications or through microbenchmarks. The authors of (CPU-and-GPU-initiated-Communication-Strategies-for-CG) implemented standard and pipelined Conjugate Gradient (CG) using MPI, NCCL/RCCL, and NVSHMEM on both NVIDIA and AMD GPUs. They found that avoiding CPU-GPU synchronization via streams improved CG performance significantly by 5–15%. For NVIDIA systems, they suggested using NCCL for small-message AllReduce but using MPI for point-to-point. On AMD GPUs, MPI outperformed vendor solutions, which was attributed to the less mature vendor software.
The benefits of NVSHMEM’s GPU-initiated communication have been demonstrated in molecular dynamics simulations using GROMACS (Doijade_2025), yielding performance improvements of up to 2x over MPI. These gains are attributed to the suitability of the PGAS model, which removes the CPU from the critical path and enables implementation-level optimizations such as kernel fusion. However, with the recent introduction of the NCCL Device API, new studies suggest that its performance is now comparable to NVSHMEM for both point-to-point and all-to-all communication.
Similar advantages have been observed in graph processing applications, specifically Breadth-First Search (BFS) (10.1007/978-3-319-73814-7_6). In these workloads, the inherent irregular and fine-grained access patterns align more effectively with the NVSHMEM programming model than with standard MPI. Similar findings have been reported for other irregular workloads such as updates of distributed Bloom filters and the search of connected components in large-scale graphs (10.1145/3708035.3736073).
Other works (gpugpuinterconnect) compare NCCL (and RCCL) to GPU-Aware MPI on both collective and point-to-point operations on three supercomputers. The results show a tendency for MPI to perform better in point-to-point operations, while *CCL performs better in collectives. However, this also depends on the specific system and optimizations provided by those libraries, and sometimes the best solution depends both on the number of nodes and vector size. Moreover, the paper reports instability in RCCL at high node count, as also highlighted in a comparison between Cray MPICH and RCCL performance on the Frontier supercomputer (singh2025bigsendoffhighperformance).
Another limitation reported for both Open MPI (gpugpuinterconnect) and Cray MPICH (singh2025bigsendoffhighperformance) concerns the execution of reduction-based collectives, such as MPI_Allreduce and MPI_Reduce_scatter. Unless specifically configured with accelerator-offload components, these MPI implementations often default to a host-staging protocol: data is copied to the host, reduced using the CPU, and then copied back to the device. This round-trip data movement and reliance on CPU arithmetic capabilities significantly degrade performance compared to native GPU collective libraries, which perform reductions directly on the GPU. This was also addressed in a study optimizing the MFDn nuclear configuration interaction code (11018300). The paper demonstrated that replacing the baseline MPI/OpenACC implementation with native CUDA kernels for computation and high-performance communication protocols yielded substantial gains. For large, network-bandwidth-bound problem sizes, the NCCL/CUDA approach proved most effective, achieving up to a 4.9x speedup over the baseline by eliminating the host-staging for reductions and fully leveraging GPU-native communication and parallelism.
Despite these variations in programming and performance, all three libraries are mature, with diverse implementations available; however, new GPUCCL variants as discussed in Section 4.2.2 are actively emerging from industry and research to address evolving AI application needs.
4.4.5. Library Interactions
Figure 4. Interactions between different GPU-GPU communication libraries.
Figure 4 illustrates the software stack hierarchy from the application down to the hardware. Applications typically interface directly with high-level libraries such as MPI, GPUSHMEM, or GPUCCL. Lower-level communication libraries, such as UCX or libfabric, generally serve as middleware used by these high-level frameworks rather than being accessed directly by applications. While legacy MPI implementations often ran directly over transport interfaces (e.g., libfabric or libibverbs), modern GPU-aware MPI stacks are more modular. They typically layer on top of UCX for point-to-point operations and UCC for collectives. In some cases, such as with MVAPICH, MPI may also offload collective operations directly to GPUCCL to leverage its topology-aware optimizations.
Similarly, GPUSHMEM adopts a hybrid architecture: it utilizes UCX or direct transport interfaces (libfabric/libibverbs) for low-latency point-to-point operations, while layering on top of GPUCCL to accelerate collective routines. Finally, GPUCCL typically interfaces directly with the transport layer (libfabric or libibverbs), though it can also be configured to run on top of UCX via specific plugins for enhanced portability. Last, UCC serves as a higher-level abstraction, executing collective operations by either offloading them to GPUCCL or employing its own implementations over UCX and low-level transports.